OpenStudy (anonymous):

Om six years, Kim will be at least 3 times as old as she is now. What can you conclude about her age now?

OpenStudy (anonymous):

OM?

OpenStudy (anonymous):

ommmmm i would call her age \(m\) and solve \(m+6\geq 3m\)

OpenStudy (anonymous):

subtract \(m\) to get \(6\geq 2m\) or \(6\geq m\)

OpenStudy (anonymous):

in**

OpenStudy (anonymous):

I got 3 is greater and equal to m